solve the simultaneous
equations graphically x-2y=1 & x+2y=1

Answer by rothauserc(4718)

x-2y=1 & x+2y=1
start with standard form for each equation
y = (x/2) - (1/2) - red line in graph
y = (-x/2) + (1/2) - green line in graph
the graph for the two equations is

solution is the point (1, 0), that is, x=1 and y=0